# Navigating the Latent Space Dynamics of Neural Models

### v2

@@ 3.2 Characterizing the attractors of the latent vector field
In this section we aim to characterize what the fixed point solution of Eq 3 represents. We start by considering two base cases:
Linear maps In the linear case, the encoding map $E$ and decoding $D$ are parametrized by matrices $\mathbf{W}_{1}\in\mathbb{R}^{N\times k}$ and $\mathbf{W}_{2}\in\mathbb{R}^{k\times N}$ and the only fixed point of the map corresponds to the origin if the network is bias free or to a shift of it. The rate of convergence of the iteration in Eq 3 is established by the spectrum of $\mathbf{W}_{2}^{T}\mathbf{W}_{1}$, and the iteration is equivalent to shrinking the input in the direction of the eigenvectors with associated eigenvalue $\lambda<1$. In case the eigenvectors of the trained AEs are aligned with the optimal solution given by the Principal Component Analysis (PCA) decomposition of the data $\mathbf{X}=\mathbf{\Phi}\mathbf{\Lambda}\mathbf{\Phi}^{*}$, then the latent vector field vanished as the mapping is isometric and no contraction occurs.
Homogeneous maps In the case of homogeneous NNs, the network satisfies $F(c\mathbf{x})=c^{\alpha}F(\mathbf{x})$ with $c\in\mathbb{R}$ for some $\alpha$. For example, this holds for ReLU networks without biases with $\alpha=1$, which learn a piecewise linear mapping. The input-output mapping can be rewritten as:
$F(\mathbf{x})=J_{F}(\mathbf{x})\mathbf{x}$ (6)
A similar observation was made in for denoising networks in [24], we remark here its generality.
This equality implies that we can rewrite Equation 3 as :
Proposition 3.2.
The error $e_{t}=\|\mathbf{z}_{t}-\mathbf{z}^{*}\|$ converge exponentially to an $\epsilon$ depending on the spectral norm $\rho(J_{f}(\mathbf{z}^{*}))$, according to the formula $\frac{\log(\frac{\epsilon}{\|\mathbf{e}_{0}\|})}{\log(\|\rho(J_{f}(\mathbf{z}^{*}))\|)}$ that provides an estimate for the number of iterations $T$ to converge to the attractor.
In Figure 11 in the Appendix, we measure how well the convergence formula predicts the measured number of iterations to converge to an attractor, on a fully non linear AEs with biases, where the assumptions of the current section are less likely to hold. The error in the estimate will be higher, when the initial condition $\mathbf{z}_{0}$ is far from the attractor, as higher order terms dominate the dynamics, and the first order Taylor approximation accumulates more error.
Extreme overparametrization case When the capacity of a network exceeds the number of training examples by far, AEs enter in a overfitting regime which leads to data memorization [56, 42, 22, 24]. The network in this case learns a constant function or an approximation thereof, which can be retrieved through the iterations in Eq:3.
Between memorization and generalization
@@
Figure 2: Memorization vs Generalization: Attractors memorize the training data as a function of the rank of $J_{f}(\mathbf{z})$ by adjusting the bottleneck dimension $k$ (Left); which is inversely proportional to the amount of generalization attained by the model (center); On the right we show example of attractors transition from a strong memorization model (First row), to good generalization (Last row).
Our goal is to characterize the properties of the latent vector field and the learned attractors in the general setting. We argue that different models can perfectly fit training points reaching similar low loss, but interpolate differently outside the training support, depending on the regularization strength and the amount of overparametrization. In this context, which information do $\mathbf{z}^{*}$ store in general? We argue that neural models fall in a spectrum between memorization and generalization, depending on the strength of the regularization term in 1, and we hypothesize that attractors fully characterize where in the spectrum a model is. We consider attractors as representations summarizing the information stored in the weights of the network. This interpretation is in line with the convergences to modes of Theorem 3.1 and can be easily seen in the case of memorization. Specifically, we assume that attractors form a dictionary of $N$ signals $\mathcal{D}^{*}=\{\mathbf{z}_{1}^{*},...,\mathbf{z}_{P}^{*}\}$ from a set of initial conditions sampled from an input distribution $P_{0}$. The dictionary corresponds to, possibly nonlinear, interpolations of the training data X, therefore $\mathcal{D}^{*}=\Phi(X)$. In the memorization regime, $\Phi$ corresponds to the standard basis up to permutations, and the network approximates a constant function which can be retrieved through the iterations of Equation 3. Transitioning to a generalization regime $\Phi$ can become arbitrarily complex. We demonstrate the interplay between memorization and generalization empirically in the experimental sections below. In Figure 2, we show how a model transitions from memorization to generalization according to the strength of the regularization on $J_{f}(\mathbf{z}^{*})$, by measuring how many training examples are stored as attractors. We remark that this memorization corresponds to an over regularized regime, as opposed to the overfitting regime of extremely overparametrized networks.
@@
Setting We trained 30 convolutional AEs models on the CIFAR, MNIST, and FashionMNIST datasets, varying the bottleneck dimension $k$ from 2 to 512. This acts as a hard regularizer on $J_{f}(\mathbf{x}^{*})$, by constraining its rank to be $<=k$. We compute attractors $\mathbf{Z}^{*}$ from elements of the training set $\mathcal{X}_{train}$ by iterating $\mathbf{z}$ till convergence. We measure the degree of memorization by defining a memorization coefficient, which is given by the cosine similarity of each decoded attractor $\mathbf{x}^{*}=D(\mathbf{z}^{*})$ to its closest point in the training set, $mem(\mathbf{z}^{*})=\min_{x\in\mathcal{X}_{train}}cos(D(\mathbf{z}^{*}),\mathbf{x})$ and we report the mean and standard deviation over attractors. We measure generalization by simply reporting the error on the test set. In Figure 12 in the Appendix, we also report the rank explaining $90\%$ of the variance of the matrix of decoded attractors $\mathbf{X}^{*}$, showing that attractors corresponding to good generalization models span more directions in the input space. For additional information on the model, hyperparameters, and settings, we refer to the Appendix.
@@
Analysis of results In Figure 2 on the left, we observe that the more regularized networks ($k$ from 2 to 16) tend to memorize data, trading off generalization performance (middle plot). We remark that this kind of memorization is not due to overfitting, as was shown in [24] for diffusion models, but it happens in the underfit regime, due to the strong regularization constraint on $J_{f}(\mathbf{x})$. We show in the Appendix results in the overfitting regime, by training the models with different sample sizes and observing a similar pattern.
@@
Takeaway: Attractors capture the interplay between generalization and memorization of neural models, which corresponds to the trade-off between the reconstruction performance and regularization term of the AEs model.
@@
(a) Evolution of a 2D Latent Vector field over training iterations
(b) Memorization to generalization
(c) Number of attractors
@@
Figure 3: The dynamics of the latent vector field (a) The 2D vector field expands from a single attractor, eventually stabilizing and over-fitting because of capacity limits. Bottom: Evolution of larger capacity AEs ($k=128)$ across training. (b) Throughout training, the network first memorizes the data with a high memorization coefficient (in blue) and then generalizes, achieving a low test error (red). (c): Evolution of attractor count for training (blue), test (red), and noise (yellow) samples; (d) Attractors computed from training and from gaussian noise converge during training (green), while the separability of the trajectories (measured as FPR95, the lower the better) increase (purple).
@@
In Figure 3 we show a similar transition from a memorization regime to generalization across training. Setting We monitor the latent vector field and attractors statistics across the training dynamics of a convolutional AEs trained on the MNIST with bottleneck dimension $k=128$.
@@
Analysis of results In Figure 3(b) we show the transition from memorization, occurring at the first epochs of training, to generalization, by plotting the memorization coefficient and the test error across training, observing a trade off between the two. In the central plot, we plot the fraction of distinct attractors, computed from 5000 random elements respectively from the $P_{train},P_{test}$ and $\mathcal{N}(0,I)$, where we consider two attractors equal if $cos(\mathbf{z}_{1}^{*},\mathbf{z}_{2}^{*})>0.99$. We observe that the number of attractors grows across epochs starting from a single attractor at initialization, as we can also observe in the 2d case in Figure 3(a), with the training and test number of attractors stabilizing according to the test loss and the noise ones converging later. On the right, we plot the similarity of attractors computed from the training and the noise distributions, measured by Chamfer symmetric similarity, observing that it increases across training, as the two distributions of attractors match. Despite being able to compute similar attractors starting from noise or training points, the trajectories in latent vector to attractors are very different, according to the source distribution considered: we show this by reporting FPR95 scores for separating trajectories from noise distribution to train distribution, based on the following measure $score(\mathbf{z})=\frac{1}{N}\sum_{z_{i}\in\pi(\mathbf{z})}d(\mathbf{z}_{i},\mathbf{Z}_{train}^{*})$ where $\pi(\mathbf{z})=[\mathbf{z}_{0},...,\mathbf{z}_{N}]$ is the trajectory. FPR drastically decreases with training, demonstrating that the networks achieves separability of the two distributions of trajectories.
@@
Takeaway: As the latent vector field evolves during training, the model transitions from memorization to generalization, forming similar attractors from different input distributions, while retaining information of the source distribution in the latent trajectories.
@@ 4 Experiments on vision foundation models
In this section we demonstrate the existence and applicability of the latent dynamics on vision foundation models, including AE backbone of the Stable Diffusion model[46] variants and vision transformer masked AEs [15], showing; (i) how information stored in the weights of a pretrained model can be recovered by computing the attractors from noise (ii) how trajectories in the latent dynamics are informative to characterize the learned distribution and detect distribution shifts.
@@ 4.1 Data free weight probing
Figure 4: Data-free weight information probing of Stable Diffusion model: We plot the error (MSE) vs sparsity (number of atoms) used to reconstruct samples from diverse dataset respectively from (i) an orthonormal random basis of the latent space (blue); (ii) attractors computed from gaussian noise (red), showing that attractors consistently reconstructs samples better on all datasets. (Right) Reconstructions using $5\%$ of the atoms on ImageNet
Setting. In the following experiment, our goal is to analyze to what degree we can recover information stored in the weights of a neural model from attractors computed from Gaussian noise, without relying on any input data. We focus on the analysis of the AE of stable diffusion [46] pretrained on the large scale Laion2B dataset [47]. We sample input $\mathbf{Z}_{n}\sim\mathcal{N}(\mathbf{0},I)$ and compute the attractors corresponding to the solution of the equation $f(\mathbf{Z}_{n}^{*})=\mathbf{Z}_{n}^{*}$. Specifically, we sample $N=4096$ points which correspond to the latent dimensionality of the model $k$. We evaluate the reconstruction performance on 6 different and diverse datasets Laion2B [47],Imagenet [9], EuroSAT [18], CIFAR100 [27],PatchCamelyon [30], Places365 [31], showcasing highly different distributions of images from general domains, to medical and satellite data. We evaluate on 500 test samples from each dataset, sampled randomly, using Orthogonal Matching Pursuit (OMP) [33] to reconstruct the sample from the attractors in the latent space, as a function of the sparsity imposed by the number of atoms selected. We compare with the reconstructions from the initial orthogonal basis $Z_{0}^{*}$, which, when using all samples, span completely $\mathcal{Z}$. The reconstruction performance is measured in the output space, decoding the reconstructed latents via $D$.
[…]
@@ 4.2 Latent trajectories characterize the learned distribution
In the following experiment, we test the hypothesis that trajectories are informative on the distribution learned by the model, testing how much Theorem 3.1 holds in real cases. In particular, we want to assess how much the paths in the vector field can detect distribution shifts in the data. For a sample to be classified as out-of-distribution (OOD) we identify two relevant questions to investigate: (i) does the sample iteration converge to one of the attractors of the training data, i.e., does it share the same basin of attraction? (ii) If so, how fast does it converge to the attractor? An OOD sample can, in fact, lie inside one of the attractor basins shared with training data. In order to assess this, we propose to track the distribution of distances to attractors point along the encoding-decoding iteration of a test sample. In the former case, we expect that an OOD sample converges faster, while in the latter case, the distance term to the attractors will dominate the score.
[…]
Analysis of results. In Table we report OOD detection performance by using the following score function: for a test sample $\mathbf{z}_{test}$ we compute its trajectory $\pi_{\mathbf{z}_{test}}=[\mathbf{z}_{0},...,\mathbf{z}_{N}^{*}]$ towards attractors and compute the distance of $\pi_{\mathbf{Z}_{test}}$ w.r.t. to the set of training attractors $\mathbf{Z}_{train}^{*}$, i.e. $score(\mathbf{z})=d(\pi_{\mathbf{z}_{test}},\mathbf{Z}_{train}^{*})$. As a distance function, we employ Euclidean distance, and we aggregate the score computing the mean distance over the trajectory $score(\mathbf{z})=\frac{1}{N}\sum_{i}d(\mathbf{z}_{i},\mathbf{Z}_{train}^{*})$. We compare with a $K$-Nearest neighbor baseline, where the score for a test sample is obtained by taking the mean distance over the $K$-NN on the training dataset, where $K=2000$ in the experiments. The score proposed demonstrates how informative the latent vector field is on the training distribution. In Figure 5 on the right, we show histograms of scores for the distance to attractors and the KNN, showing again that the former method is able to tell apart in-distribution and out-of-distribution data correctly.
@@
Figure 5: Trajectories in the latent vector field characterize distribution shifts We measure out-of-distribution detection performance on ViTMAE: On the left we report scores for 4 different datasets, highly outperforming the KNN baseline. On the right histograms of scores the INaturalist dataset demonstrating much better separability between in-distribution and out-of-distribution.

### v3

@@ 3.2 Characterizing the attractors of the latent vector field
In this section, we aim to characterize what the fixed point solutions of Eq. 3 represent. In Appendix A.4 we consider the two cases of linear and homogeneous networks as examples, to build intuition on the characterization of attractors and latent vector field.
Between memorization and generalization.
@@
Figure 2: Memorization vs Generalization. Attractors memorize the training data as a function of the rank of $J_{f}(\mathbf{z})$ by adjusting the bottleneck dimension $k$ (left) which is inversely proportional to the amount of generalization attained by the model (center); On the right we show example of attractors transitioning from a strong memorization model (first row) to good generalization (last row).
Our goal is to characterize the properties of the latent vector field and the learned attractors in the general setting. We argue that different models can fit training points reaching similar low loss, but interpolate differently outside the training support, depending on the regularization strength and the amount of overparametrization.
In this context, a key question arises: what information do the attractors $\mathbf{z}^{*}$ encode? We argue that neural models lie in a spectrum between memorization and generalization, depending on the strength of the regularization term in Eq. 1. We hypothesize that attractors fully characterize where a model falls on this spectrum. To do so, we first connect attractors to the notion of generalization via the following proposition:
Proposition 3.2 (Informal, proof in Appendix A.3).
Let $\mathbf{Z}^{*}$ be a dictionary of attractors of $f=E\circ D$ in a neighborood $\Omega$ of the latent space, and let $\Pi(\mathbf{z})$ denote the projection onto the nearest attractor to a latent code $\mathbf{z}=E(\mathbf{x})$. If $D$ is $L_{D}$-Lipschitz on $\Omega$, then for any test point $\mathbf{x}$:
$\|{\mathbf{x}-F(\mathbf{x})}\|_{2}^{2}\;\leq\;\underbrace{\|{\mathbf{x}-D(\Pi(E(\mathbf{x})))}\|_{2}^{2}}_{\text{error to prototype}}+\underbrace{L_{D}^{2}\,\|E(\mathbf{x})-\Pi(E(\mathbf{x})))\|_{2}^{2}}_{\text{coverage error}}.$
In short, attractors define a dictionary for the data: when they coincide with training points, prototype error on these points vanishes but coverage is narrow (memorization regime), whereas generalization requires attractors that both cover the latent space and serve as good prototypes for unseen data.
We consider attractors as representations summarizing the information stored in the weights of the network. This interpretation is in line with the convergence of paths in the vector field to modes of the learned distribution of Theorem 1, and can be easily seen in the case of memorization.
Extreme overparametrization case. When the capacity of a network exceeds the number of training examples by far, AEs enter an overfitting regime leading to data memorization (Zhang et al., 2019; Radhakrishnan et al., 2020; Jiang and Pehlevan, 2020; Kadkhodaie et al., 2023). The network in this case learns a constant function or an approximation thereof, which can be retrieved through the iterations of Eq. 3.
3.2.1 The role of regularization
We empirically demonstrate that the balance between memorization and generalization is governed by the strength of regularization. In Figure 2, we show how increasing regularization on the Jacobian $J_{f}(\mathbf{z}^{*})$ drives the model from a memorization regime, where many training examples are stored as attractors, toward generalization. We remark that this memorization arises in an over-regularized regime, as opposed to the overfitting regime of extremely overparametrized networks.
@@ 3.2.1 The role of regularization
Setting. We trained 30 convolutional AEs on the CIFAR, MNIST, and FashionMNIST datasets, varying the bottleneck dimension $k$ from 2 to 512. This acts as a hard regularizer on $J_{f}(\mathbf{z}^{*})$, by constraining its rank to be $\leq k$. We compute attractors $\mathbf{Z}^{*}$ from elements of the training set $\mathcal{X}_{train}$ by iterating $f$ till convergence. We measure the degree of memorization by defining a memorization coefficient, which is given by the cosine similarity of each decoded attractor $\mathbf{x}^{*}=D(\mathbf{z}^{*})$ to its closest point in the training set, $\mathrm{mem}(\mathbf{z}^{*})=\min_{x\in\mathcal{X}_{train}}\cos(D(\mathbf{z}^{*}),\mathbf{x})$ and we report the mean and standard deviation over attractors. We measure generalization by simply reporting the error on the test set. In Figure 12 in the Appendix, we also report the rank explaining $90\%$ of the variance of the matrix of decoded attractors $\mathbf{X}^{*}$, showing that attractors corresponding to good generalization models span more directions in the input space. For additional information on the model, hyperparameters, and settings, we refer to Appendix D.
@@
Analysis of results. In Figure 2 on the left, we observe that the more regularized networks ($k$ from 2 to 16) tend to memorize data, trading off generalization performance (middle plot). We remark that this kind of memorization is not due to overfitting, as was shown in (Kadkhodaie et al., 2023) for diffusion models, but it happens in the underfit regime, due to the strong regularization constraint on $J_{f}(\mathbf{x})$. We show in Figure 13 in the Appendix results in the overfitting regime, by training the models with different sample sizes and observing a similar pattern.
@@
Takeaway. Attractors capture the interplay between generalization and memorization of neural models, which corresponds to the trade-off between the reconstruction performance and regularization term of the AE model.
@@
(a) Evolution of a 2D latent vector field over training iterations.
(b) Memorization to generalization.
(c) Number of attractors.
@@
Figure 3: Latent vector field dynamics. (a) The 2D vector field ($k=2$) expands from a single attractor, eventually stabilizing and over-fitting because of capacity limits. Bottom: Evolution of larger capacity AEs ($k=128)$ across training. (b) Throughout training, the network first memorizes the data with a high memorization coefficient (in blue) and then generalizes, achieving a low test error (red). (c): Evolution of attractor count for training (blue), test (red), and noise (yellow) samples; (d) Attractors computed from training and from gaussian noise converge during training (green), while the separability of the trajectories (measured as FPR95, the lower the better) increase (purple).
3.2.2 Memorization and generalization across training
In the experiment in Figure 3 we show that a similar transition from a memorization regime to generalization occurs across training.
@@ 3.2.2 Memorization and generalization across training
Setting. We monitor the latent vector field and attractors statistics across the training dynamics of a convolutional AE trained on MNIST with bottleneck dimension $k=128$ (bottom row plots in the Figure). To show qualitatively the evolution of a latent vector field, we also plot it across training epochs for AEs with bottleneck dimension $k=2$ in Figure 3(a).
@@
Analysis of results. In Figure 3(b) we show the transition from memorization, occurring at the first epochs of training, to generalization, by plotting the memorization coefficient and the test error across training, observing a trade-off between the two. In the center plot, we plot the fraction of distinct attractors, computed from 5000 random elements respectively from the $P_{train},P_{test}$ and $\mathcal{N}(0,I)$, where we consider two attractors equal if $\cos(\mathbf{z}_{1}^{*},\mathbf{z}_{2}^{*})>0.99$. Initially, the model converges to a single attractor (as also seen in the 2D example in Figure 3(a)). Over time, the number of attractors increases, stabilizing for training and test data in tandem with the test loss, while attractors from noise inputs converge more slowly. The right plot shows the Chamfer symmetric similarity between attractors from the training and noise distributions, which increases over training as the two distributions of attractors match.
Importantly, while attractors derived from noise and training data become increasingly similar over the course of training, the trajectories leading to them differ significantly. To quantify this, we compute FPR95 scores for distinguishing trajectories originating from noise versus those from training data. We define the trajectory score as $\mathrm{score}(\mathbf{z})=\frac{1}{N}\sum_{\mathbf{z}_{i}\in\pi(\mathbf{z})}d(\mathbf{z}_{i},\mathbf{Z}_{train}^{*})$ where $\pi(\mathbf{z})=[\mathbf{z}_{0},...,\mathbf{z}_{N}]$ is the trajectory. The FPR95 decreases sharply during training, showing that the network learns to separate the two types of trajectories.
@@
Takeaway. As the latent vector field evolves during training, the model transitions from memorization to generalization, forming similar attractors from different input distributions, while retaining information of the source distribution in the latent trajectories.
@@ 4 Experiments on vision foundation models
In this section we demonstrate the existence and applicability of the latent dynamics on vision foundation models, including the AE backbones of the Stable Diffusion model (Rombach et al., 2022) and vision transformer masked AEs (He et al., 2022), showing; (i) how information stored in the weights of a pretrained model can be recovered by computing the attractors from noise (ii) how trajectories in the latent dynamics are informative to characterize the learned distribution and detect distribution shifts.
@@ 4.1 Data free weight probing
Figure 4: Data-free weight information probing of Stable Diffusion model. We plot the error (MSE) vs sparsity (number of atoms) used to reconstruct samples from diverse dataset respectively from (i) an orthonormal random basis of the latent space (blue); (ii) attractors computed from gaussian noise (red), showing that attractors consistently reconstructs samples better on all datasets. (Right) Reconstructions using $5\%$ of the atoms on ImageNet
Setting. In this experiment, we investigate how much information stored in a neural network’s weights can be recovered purely from attractors computed on Gaussian noise, without access to any input data. We focus on AE component of Stable Diffusion (Rombach et al., 2022) pretrained on the large scale Laion2B dataset (Schuhmann et al., 2022).
We sample $\mathbf{Z}_{n}\sim\mathcal{N}(\mathbf{0},I)$ and compute the corresponding attractors as solutions to $f(\mathbf{Z}_{n}^{*})=\mathbf{Z}_{n}^{*}$. We generate $N=4096$ such points, matching the latent dimensionality $k$ of the model. We evaluate the reconstruction performance on 6 diverse datasets Laion2B (Schuhmann et al., 2022), Imagenet (Deng et al., 2009), EuroSAT (Helber et al., 2019), CIFAR100 (Krizhevsky et al., 2009), PatchCamelyon (Litjens et al., 2018), and Places365 (Zhou and Paffenroth, 2017). These datasets span general, medical, and satellite image domains.
For each dataset, we randomly sample 500 test examples and reconstruct them in latent space using Orthogonal Matching Pursuit (OMP) (Mallat and Zhang, 1993), with varying sparsity levels (i.e., number of atoms used). Reconstructions are decoded via $D$, and performance is compared against a baseline reconstruction using the initial orthogonal samples $\mathbf{Z}_{0}^{*}$, which, by design, fully span the latent space $\mathcal{Z}$.
Analysis of results. In Figure 4, we plot the error as a function of the number of elements (atoms) chosen to reconstruct the test samples using OMP. We compare by building a random orthogonal basis sampled in the latent space. For the latter, applying OMP corresponds to performing Principal Component Analysis (PCA) on the chosen basis. On all considered datasets, noise attractors recover test samples with lower reconstruction error, representing a better dictionary of signals. In Figure 4(a), we show qualitative reconstruction using only 5 $\%$ of the atoms. In Figures  789 in the Appendix, we include additional qualitative evidence of this phenomenon, visualizing reconstructions of random samples of the datasets as a function of the number of atoms used. In Appendix B, we report additional results on variants of the AE of different sizes.
@@ 4.2 Latent trajectories characterize the learned distribution
In the following experiment, we test the hypothesis that trajectories are informative on the distribution learned by the model, testing how well Theorem 1 holds in practice. Our goal is to evaluate whether these trajectories can be used to detect distribution shifts in the input data. To classify a sample as out-of-distribution (OOD), we focus on two key questions: (i) does the sample trajectory converge to one of the attractors of the training data, i.e., does it share the same basin of attraction? (ii) If so, how fast does it converge?
Notably, an OOD sample may still lie within a shared attractor basin. To capture both scenarios, we track the distance from each point along a test sample’s trajectory to the nearest attractor from the training set. In the former case, we expect that an OOD sample converges faster, while in the latter case, the distance term to the attractors will dominate the score.
@@
Figure 5: Trajectories in the latent vector field characterize distribution shifts We measure out-of-distribution detection performance on ViTMAE: On the left we report scores for 4 different datasets, highly outperforming the KNN baseline. On the right, histograms of scores on the INaturalist dataset, demonstrating much better separability between in-distribution and out-of-distribution.
Setting. We test this hypothesis using the ViT-MAE (He et al., 2022) architecture pretrained on ImageNet. We sample 2000 training images and compute their attractors, stopping when convergence reaches a tolerance of $10^{-5}$ or a maximum number of iterations. We test on samples from SUN397 (Xiao et al., 2016), Places365 (Zhou and Paffenroth, 2017), Texture (Cimpoi et al., 2014), and iNaturalist (Van Horn et al., 2018), standard benchmarks for OOD detection (Yang et al., 2024). We report two metrics: False Positive Rate at $95\%$ True Positive Rate (FPR95), and Area Under the Receiver Operating Characteristic Curve (AUROC). FPR95 measures what percentage of OOD data we falsely classify as ID where our threshold includes $95\%$ of ID data.
Analysis of results. In Figure 5 we report OOD detection performance by using the following score function: for a test sample $\mathbf{z}_{test}$ we compute its trajectory $\pi_{\mathbf{z}_{test}}=[\mathbf{z}_{0},...,\mathbf{z}_{N}^{*}]$ towards attractors and compute the distance of $\pi_{\mathbf{Z}_{test}}$ to the set of training attractors $\mathbf{Z}_{train}^{*}$, i.e. $\mathrm{score}(\mathbf{z})=d(\pi_{\mathbf{z}_{test}},\mathbf{Z}_{train}^{*})$. As a distance function, we employ Euclidean distance, and we aggregate the score by computing the mean distance over the trajectory $\mathrm{score}(\mathbf{z})=\frac{1}{N}\sum_{i}d(\mathbf{z}_{i},\mathbf{Z}_{train}^{*})$. We compare with a $K$-Nearest neighbor baseline, where the score for a test sample is obtained by taking the mean distance over the $K$-NN on the training dataset, where $K=2000$ in the experiments. The score proposed demonstrates how informative the latent vector field is on the training distribution. In Figure 5 on the right, we show histograms of scores for the distance to attractors and the KNN, showing again that the former method is able to tell apart in-distribution and out-of-distribution data correctly.